# CoRa: A Collision-Resistant LoRa Symbol Detector of Low Complexity

### v1

@@ I Introduction
LoRa has emerged as a particularly noteworthy solution in both industry and academia due to its flexibility, wide availability, and cost-effectiveness. LoRa can operate in unlicensed spectrum, such as the 868 MHz band in Europe, the 915 MHz band in the United States and the 2.4 GHz band worldwide. Its modulation utilizes a variant of the chirp spread spectrum (CSS) technique, which enables a broad transmission range of several kilometers in urban environments and tens of kilometers in rural areas. This is achieved through the use of a linearly modulated chirp that spreads the transmission over a wide frequency band, resulting in robustness against multipath fading, interference and Doppler effect.
LoRa transmission experience long time on air, ranging from milliseconds to several seconds. Consequently, these prolonged transmission durations increase the likelihood of frame collisions with other LoRa frames. Under certain circumstances, colliding LoRa frames can still be decoded by exploiting the capture effect, i.e., when the power of the received signal exceeds that of the collided signal by a certain threshold. In contrast, LoRa packets that collide with similar power are not decodable by standard LoRa receivers.
Recent work has studied decoding LoRa frames under collisions to increase the overall utility and bandwidth of LoRa networks and demonstrated a significant increase in network capacity. However, these approaches often entail high computational complexity, pose challenges for real-time operation or require a prior knowledge of the position of interference frames relative to the target frame. The latter is ineffective if the frame detection algorithm fails to detect interfering frames, a situation that is common under severe frame collisions or low Signal-to-Noise Ratio (SNR). These limitations challenge their implementability on real-world devices.
In this work, we propose CoRa, an alternative symbol detector designed to accurately classify LoRa symbols by employing a Bayesian classifier. CoRa is specifically engineered to identify the true symbol within a frame amidst interfering signals in the received signal, all while maintaining low computational complexity. This approach enhances the demodulation accuracy of LoRa symbols in congested environments, leading to a significant improvement in the Packet Reception Ratio (PRR). Our contributions read:
@@
    We demonstrate the ability to accurately demodulate symbols under collisions and significant channel fluctuations by employing our robust symbol detector.
@@ II-D Related work
Existing work addresses the inherent constraints of LoRa transmissions through the employ of Media Access Control (MAC) strategies. Several authors propose to employ multi-hop LoRa networks to improve energy efficiency and overall communication performance [8, 9, 10, 11, 12]. There exists a growing interest in Time-Division Multiple Access schemes for LoRa communication. Zorbas et al.  [13] propose a time-slotted scheme for LoRa traffic, which focus on applications that require frequent and very reliable communication. Haubro et al.  [14] and Alamos et al.  [15] design adaptions layer to facilitate the operation of TSCH and DSME modes, respectively, of IEEE 802.15.4 over the LoRa Physical Layer (PHY). Due to the nature of time-slotted communication, these strategies trade-off transmission delay with communication reliability. Multiple studies [16, 17, 18, 18] have examined contention-based communication in LoRa networks, targeting the facilitation of unsynchronized communication, which is a common scheme in LoRaWAN networks.
@@ III CoRa design
Figure 3: System architecture of CoRa . The blue box highlights the symbol detector block, which is the our main contribution
@@ III-A Feature extraction
We calculate two key features. The Half-Period Discriminator (HPD), which exploits wave symmetry properties and is a novel contribution of this work, and the Peak Magnitude Deviation, which measures the discrepancy from the expected peak magnitude.
@@
The Half-Period Discriminator assesses the presence of symmetric or anti-symmetric periodic frequency components within a signal.
This metric is built upon the observation that a true peak manifests as a complete waveform, exhibiting either symmetry (for even frequencies) or anti-symmetry (for odd frequencies). In contrast, interference peaks typically appear as clipped waveforms that do not conform to symmetry patterns. This distinction aids in separating the true peak from interference peaks.
@@
Equation 1 shows that the symmetry of a complete waveform $e^{j2\pi\frac{k}{N}n}$ yields $Y_{k}=0$. In contrast, the peak of a clipped waveform attains a non-zero spectrum, as the clipped signal is neither symmetric nor anti-symmetric. Figure 4 illustrates $\mathbf{X}=\left|\mathcal{F}(\mathbf{x})\right|$, the magnitude of the spectrum of a dechirped symbol under collisions (top left), along with $\mathbf{Y}$, the magnitude of the spectrum after phase inversion (top-right).
@@
As illustrated in Figure 4 (bottom-right), the HPD values at the peaks are consistent with our previous analysis. Nevertheless, noise bins also yield values between zero and one, as components around such bins may contain symmetrical or anti-symmetrical characteristics. Additionally, poorly synchronized frames, which are common in LoRa deployments, cause the true peak to manifest at a fractional frequency. As a result, the phase cancellation becomes less effective, resulting in larger HPD values. Due to these factors, the HPD feature alone is insufficient for accurately estimating the location of the true peak. This limitation motivates the calculation of an additional feature, the Peak Magnitude Deviation.
Peak Magnitude Deviation (PMD)
The Peak Magnitude Deviation (PMD) measures the difference between the observed magnitude in the FFT bin of the dechirped symbol and the expected peak magnitude. This feature enhances the peak discrimination strategy proposed by CIC [23], which merely filters out peaks that do not align with the expected peak magnitude.
where $X_{k}$ is the spectrum of the dechirped symbol at bin $k$ and $E(X_{p})$ is the expected peak magnitude. The expected value $E(|X_{p}|)$ is estimated by calculating the average magnitude of the preamble peaks, which are determined during the synchronization stage.
The magnitude of the FFT bin that encapsulates the true peak is expected to align closely with the magnitude of the preamble, resulting in PMD values approaching zero. Conversely, noise bins typically exhibit lower magnitudes than the preamble, leading to PMD values that deviate from zero. Interference peaks, on the other hand, often yield greater PMD values due two primary factors: (i) the energy of the received interference packet does not necessarily correspond to the energy of the preamble and (ii) even if the energies are similar, the misalignment of interference symbols relative to the target symbol leads to clipped waveforms, causing peaks to exhibit lower energy than that of the preamble.
Similarly to the HDP feature, the PMD feature alone is not a reliable indicator of the true peak. Instances, for which the magnitude of an interference peak aligns with the magnitude of the true peak or the calculation of the expected magnitude is compromised by noise, interference peaks may exhibit a lower PMD value than the true peak. Consequently, the PMD cannot always discriminate between the true peak and interference peaks under the conditions of symbols of comparable energy levels or of low Signal-to-Noise Ratio (SNR).
This motivates our development of a robust estimator that integrates the joint features of both HPD and PMD using a Bayesian Classifier.
@@ III-B Bayesian Classifier
Let $T_{k}\left(\mathbf{x}\right)$ and $C_{k}\left(\mathbf{x}\right)$ denotes the events that the $k$ of the spectrum of the dechirped symbol $x$ contain the true peak and the expected complete waveform, respectively. Let $h_{k}\left(\mathbf{x}\right)$ and $p_{k}\left(\mathbf{x}\right)$ be the values of the Half-Period Discriminator (HPD) and the Peak Magnitude Deviation (PMD) respectively of symbol $\mathbf{x}$ at bin $k$. Let $\mathbf{hp}_{k}\left(\mathbf{x}\right)=\left(h_{k}\left(\mathbf{x}\right),p_{k}\left(\mathbf{x}\right)\right)$. Leveraging the independence between values of consecutive data symbols, we calculate the Bayesian classifier $P\left(T_{k}\left(\mathbf{x}\right)\right)$ as the probability of containing the true peak at $k$, considering $\mathbf{x}$ as the current symbol and $\mathbf{x}^{\prime}$ as the previous symbol. Specifically, $T_{k}\left(\mathbf{x}\right)$ reflects the process whereby the symbol $\mathbf{x}$ has the true peak at $k$ and true peak in the previous symbol appears at a different location, thereby ensuring that the peak is not a preamble peak.
$P\left(T_{k}\left(\mathbf{x}\right)\right)$ $=P\left(C_{k}\left(\mathbf{x}\right)\mid\mathbf{hp}_{k}\left(\mathbf{x}\right)\cap\neg C_{k}\left(\mathbf{x}^{\prime}\right)\mid\mathbf{hp}_{k}\left(\mathbf{x}^{\prime}\right)\right)$
$=P\left(C_{k}\left(\mathbf{x}\right)\mid\mathbf{hp}_{k}\left(\mathbf{x}\right)\right)$
$\quad\cdot\left(1-P\left(C_{k}\left(\mathbf{x}^{\prime}\right)\mid\mathbf{hp}_{k}\left(\mathbf{x}^{\prime}\right)\right)\right)$ (2)
@@
Utilizing Bayes’s theorem, the posterior probability $P\left(C_{k}\mid\mathbf{hp}_{k}\left(\mathbf{x}\right)\right)$ can be estimated from the prior probability as follows:
$P\left(C_{k}\left(\mathbf{x}\right)\mid\mathbf{hp}_{k}\left(\mathbf{x}\right)\right)=\frac{P\left(\mathbf{hp}_{k}\left(\mathbf{x}\right)\mid C_{k}\left(\mathbf{x}\right)\right)\cdot P\left(C_{k}\left(\mathbf{x}\right)\right)}{P\left(\mathbf{hp}_{k}\left(\mathbf{x}\right)\right)}$ (3)
This equation indicates that by knowing the probabilities $P\left(\mathbf{hp}_{k}\left(\mathbf{x}\right)\mid C_{k}\left(\mathbf{x}\right)\right)$, $P\left(C_{k}\left(\mathbf{x}\right)\right)$ and $P\left(\mathbf{hp}_{k}\left(\mathbf{x}\right)\right)$, we can calculate the probability of the true peak using the Bayes classifier (see Equation 2).
@@ III-C Parameter estimation
The posterior probability (see Equation 3), necessary for the Bayesian classifier (see Equation 2), can be determined if $P\left(\mathbf{hp}_{k}\left(\mathbf{x}\right)\mid C_{k}\left(\mathbf{x}\right)\right)$, $P\left(C_{k}\left(\mathbf{x}\right)\right)$, and $P\left(\mathbf{hp}_{k}\left(\mathbf{x}\right)\right)$ are known. According to the law of total probability:
$P\left(\mathbf{hp}_{k}(\mathbf{x})\right)$ $=P\left(\mathbf{hp}_{k}(\mathbf{x})\mid C_{k}(\mathbf{x})\right)\cdot P\left(C_{k}(\mathbf{x})\right)$
$\quad+P\left(\mathbf{hp}_{k}(\mathbf{x})\mid\neg C_{k}(\mathbf{x})\right)\cdot P\left(\neg C_{k}(\mathbf{x})\right)$ (4)
This reduces the problem to determining $P\left(\mathbf{hp}_{k}(\mathbf{x})\mid C_{k}(\mathbf{x})\right)$, $P\left(\mathbf{hp}_{k}(\mathbf{x})\mid\neg C_{k}(\mathbf{x})\right)$, and $P\left(C_{k}(\mathbf{x})\right)$.
To estimate $P\left(\mathbf{hp}_{k}\left(\mathbf{x}\right)\mid C_{k}\right)$ and $P\left(\mathbf{hp}_{k}\left(\mathbf{x}\right)\mid\neg C_{k}\right)$, we simulate samples of $\mathbf{hp}_{k}$ at both the true peak and the interference or noise peaks. We then compute smoothed 2D histograms from these samples to calculate the discrete probabilities.
@@
To create the true peak and interference/noise sample sets, we extract $\mathbf{hp}_{k}$ samples from 100,000 dechirped symbols, each containing 256 bins. Each symbol comprises a complete waveform (which represent the true peak) and up to two interference packets, each characterized by two clipped waveforms, which are uniformly distributed in terms of their frequencies and positions. Transmission power is uniformly distributed between -15 dB and 13 dB, To account for synchronization errors, a small fractional frequency deviation, uniformly distributed within ±$\pm\frac{1}{8}$ of the Nyquist frequency, is added to the symbol frequency.
For robust estimator, we focus on challenging symbols misclassified by the baseline LoRa demodulator, where the true peak magnitude is not the most prominent. The true peak set comprises the $h_{k}$ and $p_{k}$ features of the bin containing the true peak. The interference set includes the 10 samples of $\mathbf{hp}_{k}$ feature pairs with the lowest Peak Magnitude Deviation ($p_{k}$), ensuring the Bayesian classifier emphasizes the HPD feature ($h_{k}$) for accurate classification.
Figure 5: Discrete (200x200 samples) posterior probability $P\left[C_{k}\left(\mathbf{x}\right)\mid\mathbf{hp}_{k}\left(\mathbf{x}\right)\right]$, as a function of HDP ($h_{k}$) and PMD ($p_{k}$) features, estimated from simulated data. This probability is subsequently used to compute the Bayesian classifier $P\left(T_{k}\left(\mathbf{x}\right)\right)$.
By using the expressions for the discrete probabilities and the total probability, we obtain the discrete probability $P\left[C_{k}\mid\mathbf{hp}_{k}\right]$, as shown in Figure 5.
As anticipated, low values of $h_{k}$ and $p_{k}$ , which characterize the true peak, yield high probabilities. Conversely, high values of $h_{k}$ and $p_{k}$, which are indicative of noise or interference peaks, result in low probabilities.

### v2

@@ Abstract
[…]
5pt {textblock}1(.1,0.01) If you refer to this paper, please cite the peer-reviewed publication:: J. Alamos, T. C. Schmidt, M. Wählisch.
CoRa: A Collision-Resistant LoRa Symbol Detector of Low Complexity.
Proc. of the 2025 IEEE INTERNATIONAL CONFERENCE ON COMPUTER COMMUNICATION AND NETWORKS., 2025.
@@ I Introduction
LoRa has emerged as a particularly noteworthy solution in both industry and academia due to its flexibility, wide availability, and cost-effectiveness. LoRa can operate in unlicensed spectrum, such as the 868 MHz band in Europe, the 915 MHz band in the United States and the 2.4 GHz band worldwide. It is typically used within the LoRaWAN architecture, which provides a cloud-based MAC layer to facilitate device communication. The LoRa modulation utilizes a variant of the chirp spread spectrum (CSS) technique, which enables a broad transmission range of several kilometers in urban environments and tens of kilometers in rural areas. This is achieved through the use of a linearly modulated chirp that spreads the transmission over a wide frequency band, resulting in robustness against multipath fading, interference, and the Doppler effect.
Existing studies [1] estimate 42,000 devices per square kilometer in urban deployments, with each device transmitting 100-byte payloads, 300 times per day. In the EU868 region, with 8-channel LoRaWAN gateways, assuming 35% of device employ spreading factor 7 (bandwidth 125 kHz), a gateway covering an 800 m radius would receive about 10 packets per second per channel. Given a 190 ms time-on-air (ToA) for each packet, no more than 5 packets per second can be transmitted without overlap, making collisions unavoidable. While LoRa frames can be decoded if the received signal is stronger than the colliding one (capture effect), collisions of similar power prevent decoding.
Recent work has studied the decoding of LoRa frames under collisions to increase the overall utility and bandwidth of LoRa networks and could demonstrate a significant increase in network capacity. However, these approaches often entail high computational complexity, pose challenges for real-time operation or require a prior knowledge of the position of interference frames relative to the target frame. The latter is ineffective if the frame detection algorithm fails to detect interfering frames, a situation that is common under severe frame collisions or a low Signal-to-Noise Ratio (SNR). These limitations challenge implementability on real-world devices.
In this work, we propose CoRa, an alternative symbol detector designed to accurately classify LoRa symbols by employing a Bayesian classifier. CoRa is specifically engineered to identify the true symbol within a frame amidst interferences in the received signal, all while maintaining low computational complexity. This approach enhances the demodulation accuracy of LoRa symbols in congested environments, leading to a significant improvement in the Packet Reception Ratio (PRR). Our contributions read:
@@
    We demonstrate the ability to accurately demodulate symbols under collisions and under significant channel fluctuations by employing our robust symbol detector.
@@ II-D Related Work
Existing work addresses the inherent constraints of LoRa transmissions through the employ of Media Access Control (MAC) strategies. Several authors propose to employ multi-hop LoRa networks to improve energy efficiency and overall communication performance [9, 10, 11, 12, 13]. There exists a growing interest in Time-Division Multiple Access schemes for LoRa communication. Zorbas et al.  [14] propose a time-slotted scheme for LoRa traffic, which focus on applications that require frequent and very reliable communication. Haubro et al.  [15] and Alamos et al.  [16] design adaptions layer to facilitate the operation of TSCH and DSME modes, respectively, of IEEE 802.15.4 over the LoRa Physical Layer (PHY). Due to the nature of time-slotted communication, these strategies trade-off transmission delay with communication reliability. DSME-LoRa also enabled 6LoRa [17], a full-featured IPv6 network integration into the RIOT operating system [18].
Multiple studies [19, 20, 21, 22] have examined contention-based communication in LoRa networks, targeting the facilitation of unsynchronized communication, which is a common scheme in LoRaWAN networks.
@@ III CoRa Design
Figure 3: System architecture of CoRa . The blue box highlights the symbol detector block, which is our main contribution
@@ III-A Feature Extraction
We calculate two key features. The Peak Magnitude Deviation (PMD), which measures the discrepancy from the expected peak magnitude, and the Half-Period Discriminator (HPD), which exploits wave symmetry properties and is a novel contribution of this work.
Peak Magnitude Deviation (PMD)
The Peak Magnitude Deviation (PMD) quantifies the difference between the observed FFT bin magnitude of a dechirped symbol and its expected peak magnitude. While CIC [27] introduces a peak filtering method based on magnitude alignment, our work builds on this principle by developing PMD as a distinct feature, applying it as a cost function to enhance peak identification under interference. We introduce the PMD, denoted as $p_{k}$, by calculating the normalized deviation of each $X_{k}$ from the expected peak magnitude, capping the result at 1 to avoid extreme values.
where $X_{k}$ is the dechirped symbol spectrum at bin $k$ and $E(X_{p})$ is the expected peak magnitude. $E(|X_{p}|)$ is estimated by averaging preamble peak magnitudes during synchronization.
The PMD helps assess the likelihood of a bin being the true symbol. The FFT bin containing the true peak typically aligns with the preamble peak magnitude, resulting in near-zero PMD values. Interference peaks typically produce PMD values that deviate from zero due to misalignment and clipping. Similarly, noise bins often result in PMD deviations from zero as their magnitudes are generally lower than the preamble.
While PMD serves as a useful cost function, it can misrepresent the true peak in specific scenarios: (i) At low SNR scenarios, noise can distort the true symbol energy, causing deviations that increase PMD and incorrectly penalize the true peak. (ii) When interference peaks match the preamble energy, PMD values approach zero, failing to penalize interference. To address these limitations, we introduce a second feature – the Half-Period Discriminator (HPD).
@@
The Half-Period Discriminator assesses the presence of symmetric or antisymmetric periodic frequency components within a signal.
This metric is built upon the observation that a true peak manifests as a complete waveform, exhibiting either symmetry (for even frequencies) or antisymmetry (for odd frequencies). In contrast, interference peaks typically appear as clipped waveforms that do not conform to symmetry patterns. This distinction aids in separating the true peak from interference peaks.
@@
Equation 1 shows that the symmetry of a complete waveform $e^{j2\pi\frac{k}{N}n}$ yields $Y_{k}=0$. In contrast, the peak of a clipped waveform attains a non-zero spectrum, as the clipped signal is neither symmetric nor antisymmetric. Figure 4 illustrates $\mathbf{X}=\left|\mathcal{F}(\mathbf{x})\right|$, the magnitude of the spectrum of a dechirped symbol under collisions (top left), along with $\mathbf{Y}$, the magnitude of the spectrum after phase inversion (top-right).
@@
As illustrated in Figure 4 (bottom-right), the HPD values at the peaks align with our previous analysis.
However, similar to PMD, HPD alone is not sufficient to reliably identify the true peak in all scenarios: (i) Noise may exhibit symmetric or antisymmetric components, resulting in HPD values close to zero. This is shown in Figure 4 (bottom-left). (ii) While carrier frequency offset (CFO) and sample time offset (STO) can be estimated and cancelled, poor estimation can cause the true peak to appear at a fractional frequency, reducing the effectiveness of phase cancellation and leading to higher HPD values.
The challenges associated with PMD and HPD are distinct and orthogonal. This motivates the development of a robust estimator that combines both features through a Bayesian Classifier to improve peak discrimination.
@@ III-B Bayesian Classifier
Let $T_{k}\left(\mathbf{x}\right)$ and $C_{k}\left(\mathbf{x}\right)$ denotes the events that the $k$ of the spectrum of the dechirped symbol $x$ contain the true peak and the expected complete waveform, respectively. Let $p_{k}\left(\mathbf{x}\right)$ and $h_{k}\left(\mathbf{x}\right)$ be the values of the PMD and the HPD, respectively, of symbol $\mathbf{x}$ at bin $k$. Let $\mathbf{ph}_{k}\left(\mathbf{x}\right)=\left(p_{k}\left(\mathbf{x}\right),h_{k}\left(\mathbf{x}\right)\right)$. Leveraging the independence between values of consecutive data symbols, we calculate the Bayesian classifier $P\left(T_{k}\left(\mathbf{x}\right)\right)$ as the probability of containing the true peak at $k$, considering $\mathbf{x}$ as the current symbol and $\mathbf{x}^{\prime}$ as the previous symbol. Specifically, $T_{k}\left(\mathbf{x}\right)$ reflects the process whereby the symbol $\mathbf{x}$ has the true peak at $k$ and true peak in the previous symbol appears at a different location, thereby ensuring that the peak is not a preamble peak.
$P\left(T_{k}\left(\mathbf{x}\right)\right)$ $=P\left(C_{k}\left(\mathbf{x}\right)\mid\mathbf{ph}_{k}\left(\mathbf{x}\right)\cap\neg C_{k}\left(\mathbf{x}^{\prime}\right)\mid\mathbf{ph}_{k}\left(\mathbf{x}^{\prime}\right)\right)$
$=P\left(C_{k}\left(\mathbf{x}\right)\mid\mathbf{ph}_{k}\left(\mathbf{x}\right)\right)$
$\quad\cdot\left(1-P\left(C_{k}\left(\mathbf{x}^{\prime}\right)\mid\mathbf{ph}_{k}\left(\mathbf{x}^{\prime}\right)\right)\right)$ (2)
@@
Utilizing Bayes’s theorem, the posterior probability $P\left(C_{k}\mid\mathbf{ph}_{k}\left(\mathbf{x}\right)\right)$ can be estimated from the prior probability as follows:
$P\left(C_{k}\left(\mathbf{x}\right)\mid\mathbf{ph}_{k}\left(\mathbf{x}\right)\right)=\frac{P\left(\mathbf{ph}_{k}\left(\mathbf{x}\right)\mid C_{k}\left(\mathbf{x}\right)\right)\cdot P\left(C_{k}\left(\mathbf{x}\right)\right)}{P\left(\mathbf{ph}_{k}\left(\mathbf{x}\right)\right)}$ (3)
This equation indicates that by knowing the probabilities $P\left(\mathbf{ph}_{k}\left(\mathbf{x}\right)\mid C_{k}\left(\mathbf{x}\right)\right)$, $P\left(C_{k}\left(\mathbf{x}\right)\right)$ and $P\left(\mathbf{ph}_{k}\left(\mathbf{x}\right)\right)$, we can calculate the probability of the true peak using the Bayes classifier (see Equation 2).
@@ III-C Parameter Estimation
The posterior probability (see Equation 3), necessary for the Bayesian classifier (see Equation 2), can be determined if $P\left(\mathbf{ph}_{k}\left(\mathbf{x}\right)\mid C_{k}\left(\mathbf{x}\right)\right)$, $P\left(C_{k}\left(\mathbf{x}\right)\right)$, and $P\left(\mathbf{ph}_{k}\left(\mathbf{x}\right)\right)$ are known. According to the law of total probability:
$P\left(\mathbf{ph}_{k}(\mathbf{x})\right)$ $=P\left(\mathbf{ph}_{k}(\mathbf{x})\mid C_{k}(\mathbf{x})\right)\cdot P\left(C_{k}(\mathbf{x})\right)$
$\quad+P\left(\mathbf{ph}_{k}(\mathbf{x})\mid\neg C_{k}(\mathbf{x})\right)\cdot P\left(\neg C_{k}(\mathbf{x})\right)$ (4)
This reduces the problem to determining $P\left(\mathbf{ph}_{k}(\mathbf{x})\mid C_{k}(\mathbf{x})\right)$, $P\left(\mathbf{ph}_{k}(\mathbf{x})\mid\neg C_{k}(\mathbf{x})\right)$, and $P\left(C_{k}(\mathbf{x})\right)$.
To estimate $P\left(\mathbf{ph}_{k}\left(\mathbf{x}\right)\mid C_{k}\right)$ and $P\left(\mathbf{ph}_{k}\left(\mathbf{x}\right)\mid\neg C_{k}\right)$, we simulate samples of $\mathbf{ph}_{k}$ at both the true peak and the interference or noise peaks. We then compute smoothed 2D histograms from these samples to calculate the discrete probabilities.
@@
To create the true peak and interference/noise sample sets, we extract $\mathbf{ph}_{k}$ samples from 100,000 dechirped symbols, each containing 256 bins. Each symbol comprises a complete waveform (which represent the true peak) and up to two interference packets, each characterized by two clipped waveforms, which are uniformly distributed in terms of their frequencies and positions. Transmission power is uniformly distributed between -15 dB and 13 dB, To account for synchronization errors, a small fractional frequency deviation, uniformly distributed within ±$\pm\frac{1}{8}$ of the Nyquist frequency, is added to the symbol frequency.
For robust estimation, we focus on challenging symbols misclassified by the baseline LoRa demodulator, for which the true peak magnitude is not the most prominent. The true peak set comprises the $p_{k}$ and $h_{k}$ features of the bin containing the true peak. The interference set includes the 10 samples of $\mathbf{ph}_{k}$ feature pairs with the lowest Peak Magnitude Deviation ($p_{k}$), ensuring the Bayesian classifier emphasizes the HPD feature ($h_{k}$) for accurate classification.
Figure 5: Discrete (200x200 samples) posterior probability $P\left[C_{k}\left(\mathbf{x}\right)\mid\mathbf{ph}_{k}\left(\mathbf{x}\right)\right]$, as a function of PMD ($p_{k}$) and HPD ($h_{k}$) features, estimated from simulated data. This probability is subsequently used to compute the Bayesian classifier $P\left(T_{k}\left(\mathbf{x}\right)\right)$.
By using the expressions for the discrete probabilities and the total probability, we obtain the discrete probability $P\left[C_{k}\mid\mathbf{ph}_{k}\right]$ depicted in Figure 5. As anticipated, low values of $p_{k}$ and $h_{k}$, which characterize the true peak, yield high probabilities. Conversely, high values of $p_{k}$ and $h_{k}$, which are indicative of noise or interference peaks, result in low probabilities.